OpenStudy (anonymous):

Use the table below to answer this question: x −3 0 1 y -3 3 5 Find the average rate of change for the given function from x = −3 to x = 1. a) 2 b) −3 c) −1 d) −2

OpenStudy (anonymous):

Do the x and y values correspond so the first x value is the input for the first y value?

OpenStudy (anonymous):

uhhh what

OpenStudy (anonymous):

On the original table, are the x and y values side by side?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay so to find the average rate of change, you take the final value of y and subtract the initial value of y then divide this by the final value of x minus the initial value of x: \[\frac{ (5-(-3)) }{ (1-(-3)) }=\frac{ (5+3) }{ (1+3) }=\frac{ 8 }{ 4 }=2\]

OpenStudy (anonymous):

I can use that formula for all questions like this one??

OpenStudy (anonymous):

The formula for average rate of change is: \[(y _{f}-y _{1})/(x _{f}-x _{i})\] Where the subscript f means final and the subscript i means initial. For example in this problem the initial y value is -3 and the final y value is 5.

OpenStudy (anonymous):

You can use that formula for all average rate of change problems.

OpenStudy (anonymous):

Wow, thank you so much!!